# Fast post-process Bayesian inference with Variational Sparse Bayesian Quadrature

### v3

@@ A.4 Sparse Bayesian quadrature
To compute the variational posterior of vsbq, we need to compute integrals of Gaussian distributions against the sgpr posterior with heteroskedastic noise. We provide here analytical formulae for this purpose. The following formulae are written in the zero-mean case; the non-zero mean case can be straightforwardly derived from this one (i.e., by considering $f-m$). We follow Ghahramani and Rasmussen [2002a]. Here, we denote with $\bm{\Sigma}_{\ell}$ the diagonal matrix of the parameters in the covariance kernel and with $\sigma_{f}$ the output scale, so that $\kappa(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\Lambda\mathcal{N}(\mathbf{x};\mathbf{x}^{\prime},\bm{\Sigma}_{\ell})$, with $\Lambda\equiv\sqrt{(2\pi)^{D}\left|\bm{\Sigma}_{\ell}\right|}$. [^1] This formulation of the squared exponential kernel is equivalent to the main paper, but makes it easier to apply Gaussian identities.
@@
$\begin{split}\mathbb{E}\left[\mathcal{I}_{j}\right]=&\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{j},\bm{\Sigma}_{j})f(\tilde{\mathbf{x}}\mid\textbf{u})\psi(\textbf{u})\mathrm{d}\tilde{\mathbf{x}}\mathrm{d}\textbf{u}\\ =&\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{j},\bm{\Sigma}_{j})\mu_{\bm{\psi}}(\tilde{\mathbf{x}})\mathrm{d}\tilde{\mathbf{x}},\\ =&\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{j},\bm{\Sigma}_{j})\kappa(\tilde{\mathbf{x}},\mathbf{Z})\bm{\Sigma}\mathbf{K}_{\mathbf{Z},\mathbf{X}}\mathbf{S}^{-1}\mathbf{y}\;\mathrm{d}\tilde{\mathbf{x}}\\ =&\;\bm{w}_{j}^{T}\left[\bm{\Sigma}\mathbf{K}_{\mathbf{Z},\mathbf{X}}\mathbf{S}^{-1}\right]\mathbf{y},\end{split}$
@@ Appendix B Proofs
In this section, we provide proofs for the lemmas in the main paper.
@@
Assume that $\text{KL}(p(\mathbf{f},\mathbf{u}\mid\mathbf{y})\|p(\mathbf{f},\mathbf{u}\mid\bm{\psi}))<\gamma$. Then, for any $\ell>0$ there exists $K_{\ell}$ such that, for any $\mathbf{x}^{*}$, $\|\mathbb{E}[{f}(\mathbf{x}^{*})^{\ell}]-\mathbb{E}[{f}_{e}(\mathbf{x}^{*})^{\ell}]\|<K_{\ell}\sqrt{\gamma/2}$. There also exists $K_{\mathrm{e}}$ such that, for any $\mathbf{x}^{*}$, $\|\mathbb{E}[\mathrm{exp}({f}(\mathbf{x}^{*}))]-\mathbb{E}[\mathrm{exp}({f}_{e}(\mathbf{x}^{*}))]\|<K_{\mathrm{e}}\sqrt{\gamma/2}$.
@@
$\text{KL}(p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y})\|p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi}))$
$=$ $\int p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})\log\left(\frac{p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})}{p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y})}\right)\mathrm{d}\textbf{f}\mathrm{d}\textbf{u}\mathrm{d}f^{*}$
$=$ $\int p(f^{*},\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})\log\left(\frac{p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})}{p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y})}\right)\mathrm{d}\textbf{f}\mathrm{d}\textbf{u}\mathrm{d}f^{*}$
$=$ $\int p(f^{*},\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})\log\left(\frac{p(\mathbf{y})p(\textbf{f},f^{*}\mid\textbf{u})\tilde{p}(\textbf{u})}{p(\mathbf{y}\mid\textbf{f})p(\textbf{f},f^{*}\mid\textbf{u})\tilde{p}(\textbf{u})}\right)\mathrm{d}\textbf{f}\mathrm{d}\textbf{u}\mathrm{d}f^{*}$
$=$ $\int p(\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})\log\left(\frac{p(\mathbf{y})p(\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})}{p(\mathbf{y}\mid\textbf{f})p(\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})}\right)\mathrm{d}\textbf{f}\mathrm{d}\textbf{u}$
$=$ $\text{KL}(p(\textbf{f},\textbf{u}\mid\mathbf{y})\|p(\textbf{f},\textbf{u}\mid\bm{\psi})).$
@@
$\|p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y})-p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})\|_{TV}\leq\sqrt{\text{KL}(p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y})\|p(\textbf{f},\textbf{u},f^{*})\mid\bm{\psi})/2}<\sqrt{\gamma/2},$
@@
$\left\|\int f^{*\ell}(p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi}))\mathrm{d}f^{*}\right\|\leq K_{\ell}\sqrt{\gamma/2}.$ (S2)
@@
$\left\|\int f^{*\ell}(p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi}))\mathrm{d}f^{*}\right\|$$\leq\left\|\int_{\mathbb{R}\backslash A}f^{*\ell}(p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi}))\mathrm{d}f^{*}\right\|$
$\hskip 28.45274pt+\left\|\int_{A}f^{*\ell}(p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi}))\mathrm{d}f^{*}\right\|$
$\leq K_{\ell}^{\prime\prime}\sqrt{\gamma/2}+K^{\prime}_{\ell}\int_{A}\|p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi})\|\mathrm{d}f^{*}$
$\leq K_{\ell}^{\prime\prime}\sqrt{\gamma/2}+K^{\prime}_{\ell}\int_{\mathbb{R}}\|p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi})\|\mathrm{d}f^{*}$
@@ C.1 Implementation details
In this section, we describe the implementation details of variational sparse Bayesian quadrature (vsbq), including noise shaping, choice of hyperparameters, and variational inference details.
@@
As discussed by Gammal et al. [2022], we can set a reasonable threshold in $D$ dimensions by considering a multivariate normal distribution in dimension $D$. The log density of a multivariate normal distribution is proportional to the sum of $D$ independent standard 1D Gaussian random variables. By defining $\Delta_{y}=2\left[\max(\log p)-\log p\right]$, we have $\Delta_{y}\sim\chi^{2}_{D}$. Further, the threshold for a “$n$-$\sigma$ contour” [Gammal et al., 2022] is,
@@ C.5 MAP estimates via Bayesian Adaptive Direct Search (BADS)
We hypothesize that bads can be a worse choice than cma-es for our purpose of post-process inference exactly for the reasons that make bads a better optimization algorithm on these problems [Acerbi and Ma, 2017]. Specifically, an optimization algorithm that converges to the global optimum quickly and efficiently is not ideal for post-process inference, in that a more exploratory population-based algorithm, such as cma-es, provides better coverage and more information about the shape of the posterior landscape. For this reason, we recommend cma-es over bads for the purpose of post-process inference. Future work could focus on developing an optimization algorithm that is good at both tasks, that is MAP estimation and subsequent posterior approximation.
@@ C.6 Visualization of posteriors
We visualize posterior distributions as ‘corner plots’, i.e., a plot with 1D and all pairwise 2D marginals. For visualization of individual posteriors obtained by the algorithms, for all problems, we report example solutions obtained from a run with the same random seed (see Figure S5, S9, S13, and S17).
@@ C.7 Runtime analysis
In this section, we present an analysis of the runtime for our method and baselines. Such an analysis is important for a post-process inference method, which ideally should only take a relatively short time (e.g., a few minutes).333bbvi methods are not considered post-processing inference techniques and serve only as a reference baseline for posterior approximation accuracy. Their runtime primarily depends on the total number of target function evaluations and the grid search for tuning learning rates. For vsbq and nnr, we report the wall-clock runtime based on 5 independent runs with different training data sets (i.e., traces from the MAP optimizations). For laplace, the runtime is based on 5 runs for numerically estimating the Hessian matrix. We measure the runtime on both CPU and GPU platforms. [^4] It is worth noting that, in the case of GPU runtime measurement, only the sparse GP and neural network fitting were carried out with a GPU. Due to our current implementation, the stochastic variational inference optimization part for vsbq and nnr was always conducted using the CPU. Future efforts to port the stochastic variational inference procedure completely to GPU could further enhance the computational performance of both methods. Specifically, we utilize an AMD EPYC 7452 32-core Processor for CPU computations and an NVIDIA A100 for GPU computations.

### v4

@@ Acknowledgments
Declarations
Supplementary information
The supplementary material contains mathematical proofs, implementation details, additional results, and extended explanations omitted from the main text.
Data availability
The data for benchmark problems and an implementation of our algorithm are available at https://github.com/acerbilab/vsbq.
Author contribution
Conceptualization: C.L, G.C, L.A; Methodology: C.L, G.C, L.A; Theoretical analysis: C.L, G.C; Experiment design: all authors; Experiment investigation: C.L; Writing: C.L, G.C, L.A; Review and editing: all authors; Supervision: L.A.
@@ A.4 Sparse Bayesian quadrature
To compute the variational posterior of vsbq, we need to compute integrals of Gaussian distributions against the sgpr posterior with heteroskedastic noise. We provide here analytical formulae for this purpose. The following formulae are written in the zero-mean case; the non-zero mean case can be straightforwardly derived from this one (i.e., by considering $f-m$). We follow Rasmussen and Ghahramani [2002]. Here, we denote with $\bm{\Sigma}_{\ell}$ the diagonal matrix of the parameters in the covariance kernel and with $\sigma_{f}$ the output scale, so that $\kappa(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\Lambda\mathcal{N}(\mathbf{x};\mathbf{x}^{\prime},\bm{\Sigma}_{\ell})$, with $\Lambda\equiv\sqrt{(2\pi)^{D}\left|\bm{\Sigma}_{\ell}\right|}$. [^1] This formulation of the squared exponential kernel is equivalent to the main paper, but makes it easier to apply Gaussian identities.
@@
$\begin{split}\mathbb{E}\left[\mathcal{I}_{j}\right]=&\int\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{j},\bm{\Sigma}_{j})f(\tilde{\mathbf{x}}\mid\textbf{u})\psi(\textbf{u})\mathrm{d}\tilde{\mathbf{x}}\mathrm{d}\textbf{u}\\ =&\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{j},\bm{\Sigma}_{j})\mu_{\bm{\psi}}(\tilde{\mathbf{x}})\mathrm{d}\tilde{\mathbf{x}},\\ =&\int\mathcal{N}(\tilde{\mathbf{x}};\bm{\mu}_{j},\bm{\Sigma}_{j})\kappa(\tilde{\mathbf{x}},\mathbf{Z})\bm{\Sigma}\mathbf{K}_{\mathbf{Z},\mathbf{X}}\mathbf{S}^{-1}\mathbf{y}\;\mathrm{d}\tilde{\mathbf{x}}\\ =&\;\bm{w}_{j}^{T}\left[\bm{\Sigma}\mathbf{K}_{\mathbf{Z},\mathbf{X}}\mathbf{S}^{-1}\right]\mathbf{y},\end{split}$
@@ Appendix B Proofs
In this section, we provide proofs for the lemmas in the main paper. We note $p(\mathbf{f},\mathbf{u}\mid\bm{\psi}))$ the sparse GP posterior associated with $\mathbf{f}$ and $\mathbf{u}$, given the optimal variational parameters $\bm{\psi}$.
@@
Assume that $D_{\text{KL}}(p(\mathbf{f},\mathbf{u}\mid\bm{\psi})\|p(\mathbf{f},\mathbf{u}\mid\mathbf{y}))<\gamma$. Then, for any $\ell>0$ there exists $K_{\ell}$ such that, for any $\mathbf{x}^{*}$, $|\mathbb{E}[{f}(\mathbf{x}^{*})^{\ell}]-\mathbb{E}[{f}_{e}(\mathbf{x}^{*})^{\ell}]|<K_{\ell}\sqrt{\gamma/2}$. There also exists $K_{\mathrm{e}}$ such that, for any $\mathbf{x}^{*}$, $|\mathbb{E}[\mathrm{exp}({f}(\mathbf{x}^{*}))]-\mathbb{E}[\mathrm{exp}({f}_{e}(\mathbf{x}^{*}))]|<K_{\mathrm{e}}\sqrt{\gamma/2}$.
@@
$D_{\text{KL}}(p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})\|p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y}))$
$=$ $\int\int\int p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})\log\left(\frac{p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})}{p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y})}\right)\mathrm{d}\textbf{f}\mathrm{d}\textbf{u}\mathrm{d}f^{*}$
$=$ $\int\int\int p(f^{*},\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})\log\left(\frac{p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})}{p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y})}\right)\mathrm{d}\textbf{f}\mathrm{d}\textbf{u}\mathrm{d}f^{*}$
$=$ $\int\int\int p(f^{*},\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})\log\left(\frac{p(\mathbf{y})p(\textbf{f},f^{*}\mid\textbf{u})\tilde{p}(\textbf{u})}{p(\mathbf{y}\mid\textbf{f})p(\textbf{f},f^{*}\mid\textbf{u})p(\textbf{u})}\right)\mathrm{d}\textbf{f}\mathrm{d}\textbf{u}\mathrm{d}f^{*}$
$=$ $\int\int p(\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})\log\left(\frac{p(\mathbf{y})p(\textbf{f}\mid\textbf{u})\tilde{p}(\textbf{u})}{p(\mathbf{y}\mid\textbf{f})p(\textbf{f}\mid\textbf{u})p(\textbf{u})}\right)\mathrm{d}\textbf{f}\mathrm{d}\textbf{u}$
$=$ $D_{\text{KL}}(p(\textbf{f},\textbf{u}\mid\bm{\psi})\|p(\textbf{f},\textbf{u}\mid\mathbf{y})).$
@@
$\|p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y})-p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})\|_{TV}\leq\sqrt{D_{\text{KL}}(p(\textbf{f},\textbf{u},f^{*}\mid\bm{\psi})\|p(\textbf{f},\textbf{u},f^{*}\mid\mathbf{y}))/2}<\sqrt{\gamma/2},$
@@
$\left|\int f^{*\ell}(p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi}))\mathrm{d}f^{*}\right|\leq K_{\ell}\sqrt{\gamma/2}.$ (S2)
@@
$|\displaystyle\left|\int f^{*\ell}(p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi}))\mathrm{d}f^{*}\right|$\$\leq\left|\int_{\mathbb{R}\backslash A}f^{*\ell}(p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi}))\mathrm{d}f^{*}\right|$
$\hskip 28.45274pt+\left|\int_{A}f^{*\ell}(p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi}))\mathrm{d}f^{*}\right|$
$\leq K_{\ell}^{\prime\prime}\sqrt{\gamma/2}+K^{\prime}_{\ell}\int_{A}|p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi})|\mathrm{d}f^{*}$
$\leq K_{\ell}^{\prime\prime}\sqrt{\gamma/2}+K^{\prime}_{\ell}\int_{\mathbb{R}}|p(f^{*}\mid\mathbf{y})-p(f^{*}\mid\bm{\psi})|\mathrm{d}f^{*}$
@@ C.1 Implementation details
In this section, we describe the implementation details of variational sparse Bayesian quadrature (vsbq), including noise shaping, choice of hyperparameters, and variational inference details. Algorithm 1 summarizes the complete procedure of vsbq.
Input: Evaluation traces $\left(\mathbf{X},\mathbf{y},\mathbf{s}\right)=(\mathbf{x}_{n},y_{n},s_{n})_{n=1}^{N}$ from MAP optimizations
Output: Posterior approximation $q_{\phi}$, estimated surrogate elbo mean $\overline{\textsc{elbo}}$ and its standard deviation $\textsc{elbo}_{\text{sd}}$
Step 1: Trimming of evaluations
Compute lcb$(\mathbf{x}_{n})=y_{n}-\beta s_{n}$, ucb$(\mathbf{x}_{n})=y_{n}+\beta s_{n}$ for all $n$;
Set $\textsc{lcb}_{\text{max}}=\max_{n}(\textsc{lcb}(\mathbf{x}_{n}))$;
Discard $\mathbf{x}_{n}$ where $\textsc{lcb}_{\text{max}}-\textsc{ucb}(\mathbf{x}_{n})>\eta_{\text{trim}}$;
Retain remaining evaluations as $\left(\mathbf{X},\mathbf{y},\mathbf{s}\right)$;
Step 2: Sparse GP fitting
Initialize sparse GP hyperparameters by fitting an exact GP to a stratified $K$-means subset of $\left(\mathbf{X},\mathbf{y},\mathbf{s}\right)$;
repeat
       Select $M$ inducing points $\mathbf{Z}$ via greedy variance selection (see Section 3.2 and [Burt et al., 2020]);
       Update sparse GP hyperparameters via maximizing GP-elbo in Eq. 24;
until no improvement in GP-elbo;
Compute the sparse GP posterior given observations $\left(\mathbf{X},\mathbf{y},\mathbf{s}\right)$ (see Eq. 18 and Eq. 19);
Obtain a sparse GP surrogate $f$ for the target log-joint density function $f_{0}$;
Step 3: Variational inference with sparse Bayesian quadrature
Initialize variational posterior $q_{\phi}$ as a mixture of $K$ multivariate Gaussians in Eq. 20;
repeat
       Compute analytically the expected log joint $\mathbb{E}_{f}\left[\mathbb{E}_{\bm{\phi}}\left[f\right]\right]$ and its variance $\mathrm{Var}_{f}\left[\mathbb{E}_{\bm{\phi}}\left[f\right]\right]$, via sparse Bayesian quadrature (see Eq. 22 and Eq. 23) ;
       Maximize $\overline{\textsc{elbo}}$: $\mathbb{E}_{f}\left[\mathbb{E}_{\bm{\phi}}\left[f\right]\right]+\mathcal{H}[q_{\bm{\phi}}]$ using reparameterized stochastic gradients with the Adam optimizer [Kingma and Ba, 2014]; [^3] The entropy $\mathcal{H}[q_{\phi}]$ for Gaussian mixtures lacks a closed-form; we follow Acerbi et al. [2018] for stochastic estimation.
      Update $q_{\phi}$ parameters;
until convergence of $\overline{\textsc{elbo}}$ or max iterations reached;
return $q_{\phi}$, $\overline{\textsc{elbo}}$, $\textsc{elbo}_{\text{sd}}=\sqrt{\mathrm{Var}_{f}\left[\mathbb{E}_{\bm{\phi}}\left[f\right]\right]}$
Algorithm 1 Variational Sparse Bayesian Quadrature (VSBQ)
@@
As discussed by El Gammal et al. [2023], we can set a reasonable threshold in $D$ dimensions by considering a multivariate normal distribution in dimension $D$. The log density of a multivariate normal distribution is proportional to the sum of $D$ independent standard 1D Gaussian random variables. By defining $\Delta_{y}=2\left[\max(\log p)-\log p\right]$, we have $\Delta_{y}\sim\chi^{2}_{D}$. Further, the threshold for a “$n$-$\sigma$ contour” [El Gammal et al., 2023] is,
@@ C.5 MAP estimates via Bayesian Adaptive Direct Search (BADS)
We hypothesize that bads can be a worse choice than cma-es for our purpose of post-process inference, exactly for the reasons that make bads a better optimization algorithm on these problems [Acerbi and Ma, 2017]. Specifically, an optimization algorithm that converges to the global optimum quickly and efficiently is not ideal for post-process inference, in that a more exploratory population-based algorithm, such as cma-es, provides better coverage and more information about the shape of the posterior landscape. For this reason, we recommend cma-es over bads for the purpose of post-process inference. Future work could explore strategies to augment the initial set of evaluations to improve coverage and enhance approximation quality.
@@
Furthermore, as in the main text, we study the sensitivity to the number of target evaluations when using the bads optimizer. As shown in Figure S4, in contrast to cma-es, vsbq is more sensitive to the number of evaluations under bads, likely due to poorer coverage of the posterior. Nonetheless, vsbq consistently outperforms nnr across most tested settings.
(a) Bayesian timing model ($\sigma_{\text{obs}}=3$).
(b) Multisensory causal inference model ($\sigma_{\text{obs}}=0$).
Figure S4: Sensitivity to the number of target evaluations, with BADS optimizer. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the number of target evaluations, for two benchmark problems. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML $=1$, MMTV $=0.2$, GsKL $=1/8$). Compared to nnr, vsbq achieves consistently better performance across most settings.
@@ C.6 Visualization of posteriors
We visualize posterior distributions as ‘corner plots’, i.e., a plot with 1D and all pairwise 2D marginals. For visualization of individual posteriors obtained by the algorithms, for all problems, we report example solutions obtained from a run with the same random seed (see Figure S6, S10, S14, and S18).
@@ C.7 Runtime analysis
In this section, we present an analysis of the runtime for our method and baselines. Such an analysis is important for a post-process inference method, which ideally should only take a relatively short time (e.g., a few minutes).444bbvi methods are not considered post-processing inference techniques and serve only as a reference baseline for posterior approximation accuracy. Their runtime primarily depends on the total number of target function evaluations and the grid search for tuning learning rates. For vsbq and nnr, we report the wall-clock runtime based on 5 independent runs with different training data sets (i.e., traces from the MAP optimizations). For laplace, the runtime is based on 5 runs for numerically estimating the Hessian matrix. We measure the runtime on both CPU and GPU platforms. [^5] It is worth noting that, in the case of GPU runtime measurement, only the sparse GP and neural network fitting were carried out with a GPU. Due to our current implementation, the stochastic variational inference optimization part for vsbq and nnr was always conducted using the CPU. Future efforts to port the stochastic variational inference procedure completely to GPU could further enhance the computational performance of both methods. Specifically, we utilize an AMD EPYC 7452 32-core Processor for CPU computations and an NVIDIA A100 for GPU computations.